# Perception, Reason, Think, and Plan: A Survey on Large Multimodal Reasoning Models

### v1

@@ 3.3.3 Multimodal-R1
The DPO in reinforcement learning has been widely used to enhance the reasoning capabilities of large multimodal models in recent years. RLHF-V (Yu et al., 2024a), LLaVA-Reasoner (Zhang et al., 2024e) and Insight-V (Dong et al., 2024b), by leveraging a large amount of self-constructed preference data and directly applying the DPO algorithm for training, have somewhat improved the reasoning ability of the models. MMPR (Wang et al., 2024g) made modifications to the DPO algorithm, adding quality loss obtained from a Binary Classifier and generation loss from traditional SFT on top of the DPO Preference loss, which effectively enhanced the model’s CoT capabilities.
With the success of Deepseek-R1, the GRPO algorithm began to be widely applied in multimodal large models. Works including MM-EUREKA (Meng et al., 2025),Vt-R1 (Zhou et al., 2025), LMM-R1 (Yingzhe et al., 2025), R1-V (Chen et al., 2025b) , by adopting a similar approach to the text domain, have applied the GRPO algorithm to mathematical geometry problems, successfully demonstrating the phenomenon of reflection. VLM-R1 (Shen et al., 2025), Visual-RFT (Liu et al., 2025f), and Seg-Zero (Yuqi et al., 2025) utilize the GRPO algorithm to enhance the visual capabilities of multimodal large language models, such as grounding, detection, and classification. This reinforcement learning approach has successfully led to improvements in the model’s visual capabilities. Besides, works includes Video-R1 (Feng et al., 2025b) and VideoChat-R1 (Li et al., 2025g) have introduced the GRPO algorithm into the video modality, while R1-Omni (Zhao et al., 2025c) has further extended it to the audio modality. Despite this, existing work is often limited to specific tasks, and current multimodal large models have not yet been able to generalize the long-chain-of-thought abilities learned from tasks such as mathematics to the model’s general capabilities, as seen with Deepseek-R1.
@@ 4.1 Experimental Findings
EmbodiedBench (Yang et al., 2025b) Multimodal Understanding, Spatial Reasoning T, I High and Low Action Levels
EmbodiedEval (Cheng et al., 2025) Attribute QA, Spatial Reasoning T, I Broad Abilities Assessment
@@
BrowseComp (Wei et al., 2025a) Web Browsing T Easy to Verify but Hard to Solve
SWE-Bench Multimodal (Yang et al., 2024b) Code T, I Image Included in Problem Statement
@@
OmniMMI (Wang et al., 2025g) VQA, Proactive Reasoning T, V, A In Streaming Video Context
@@
Recent studies have introduced a series of omni-modal benchmarks designed to evaluate the ability of LMRMs to perform unified understanding and reasoning across various data types (e.g. images, audio, text, and video). For example, OmniMMI (Wang et al., 2025g) aims to comprehensively assess the interactive capabilities of streaming video contexts in open-world environments. Experimental results reveal that even commercial models, such as Gemini-1.5-Pro and GPT-4o, achieve an average accuracy of less than 20%. When tasks require unified modality understanding (OmniBench (Li et al., 2024j), TaskAnything and JudgeAnything (Pu et al., 2025), MixEvalL-X (Ni et al., 2024b)), the performance of both open-source and closed-source models is significantly lower than under single-modal conditions. Specifically, in the Audio-Video Question Answering (AVQA) task, such as WorldSense (Hong et al., 2025), Claude 3.5 Sonnet only achieves an average accuracy of 35%, while the best-performing open-source model only achieves an accuracy of 25%. In the case of more challenging multimodal reasoning tasks, such as BabelBench (Wang et al., 2024i) and OmnixR (Chen et al., 2024e), the performance of all models declines sharply as the number of modalities increases. This suggests that models struggle to generate reasoning paths for image, video, and audio inputs compared to text inputs. These findings collectively highlight that current LMRMs are not yet capable of effectively processing omni-modal inputs.
@@
A diverse range of tasks highlights the complexity and breadth of multimodal agent evaluation settings. These include AgentBench’s multi-environment tasks (Liu et al., 2023b, 2024d), WorFBench’s intricate workflow planning scenarios (Qiao et al., 2024), OSWorld’s and AndroidWorld’s full operating system interactions (Xie et al., 2024a; Rawles et al., 2024), EmbodiedBench’s vision-based navigation and manipulation challenges (Yang et al., 2025b), VisualWebArena’s visually grounded web tasks (Koh et al., 2024), and GAIA’s open-ended, tool-augmented queries (Hu et al., 2023). Together, these benchmarks span a wide spectrum of task types and modalities (e.g., text and vision), encompassing both realistic and tool-augmented environments.
Regarding the performance of LMRMs on agent benchmarks, these models generally lead current performance and have made notable progress (Team, 2024, 2025a; Yao et al., 2024b). However, even state-of-the-art models consistently fall short of human-level reliability and struggle with complex, open-ended tasks. Evaluations across benchmarks repeatedly expose common bottlenecks: models often fail at real-world grounding (Gou et al., 2025; Zheng et al., 2024a), coherent long-horizon reasoning and planning (Qian et al., 2025), seamless integration with external tools (Wang et al., 2025d), and maintaining robustness across diverse modalities and domains (Chu et al., 2025). For example, in the BrowseComp benchmark (Wei et al., 2025a), GPT-4o achieves only 0.6% accuracy, rising to 1.9% with browsing tools, highlighting weak tool-interactive planning capability. OpenAI’s reasoning model o1 reaches 9.9%, but still leaves significant room for improvement. Notably, OpenAI Deep Research, with targeted tuning for web search, completes 51.5% of tasks via autonomous iterative tool calling and reasoning. The experimental results highlight that current large reasoning models remain deficient in long-horizon reasoning and adaptive planning, which may require specific tuning and architectural enhancements to evolve into truly native agentic systems.
@@
Recently, OpenAI released o3 and o4-mini, providing full agentic access to ChatGPT tools and enabling models to "think with images." The integration of visual content enhances multimodal reasoning directly within the thought process. For example, in Figure 6, o3 demonstrates a clear task decomposition during an 8 minute and 13 second thought process. It effectively determines the best way to crop each sub-figure through trial and error, ultimately arriving at the correct solution.
Beyond visual reasoning, we evaluated o3’s capabilities in file processing, puzzle solving, location identification, and multimedia content creation. As illustrated in Figure 7 and 8, o3 exhibits strong performance in complex multimodal problem-solving by capturing and leveraging subtle clues in images. However, several challenges are identified: 1) Language knowledge can interfere with visual input. As the finger counting case shown n Figure 8, o3 mistakenly identifies the image as the standard raised hand emoji showing four fingers plus a thumb, despite the image clearly displaying six fingers. 2) OpenAI o3 struggles with input file handling and multimedia content generation. Due to tool constraints and the lack of Internet access in coding environments, file processing and multimedia creation often result in inaccuracies. In the resume information collection case in Figure 8, phone numbers parsed from resume PDFs can be incorrect, and o3 hallucinates candidates’ project experiences by reusing similar content. Additionally, in multimedia creation cases in Figure 7, the generated frames fail to adhere to the "red panda" instructions, and o3 is unable to support interleaved text-image generation. 3) OpenAI o3 may fabricate reasoning in its thought process. It occasionally "lies" about its reasoning, constructing incorrect rationales for potentially correct answers (e.g., the puzzle-solving case in Figure 7). This problem needs to be solved urgently, as it could lead to the model attempting to deceive users during the post-training process. In fact, it highlights that the model has not yet mastered the relevant thinking logic to solve the problem.
@@ 4.2 Capability of N-LMRMs
Search-o1 (Li et al., 2025e) 32B T T Training-Free Multi-Hop QA, Math Agentic Search-Augmented Reasoning
@@
Baichuan-Omni-1.5 (Li et al., 2025h) 7B T, I, A, V T, A Pretrain+SFT VQA, OCR, ASR, Math, GeneralQA Leading Medical Image Understanding
@@
R1-Omni (Zhao et al., 2025c) 0.5B T, A, V T RL Emotion Recognition RL with Verifiable Reward
Janus-Pro (Chen et al., 2025d) 1B, 7B T, I T, I Pretrain+SFT Multimodal Understanding, Text-to-Image Decoupling Visual Encoding For Understanding and Generation
AnyGPT (Zhan et al., 2024) 7B T, I, A T, I, A Pretrain Multimodal-to-Text and Text-to-Multimodal Discrete Representations For Unified Processing
Uni-MoE (Li et al., 2025j) 13B, 20B, 22B, 37B T, I, A, V T Pretrain+SFT VQA, AQA Modality-Specific Encoders with Connectors for Unified Representation
@@
Moreover, Search-o1 (Li et al., 2025e) utilizes external knowledge retrieval during the reasoning process to fill gaps in their understanding. R1-Searcher (Song et al., 2025b) and DeepResearcher (Zheng et al., 2025e) enhance their ability to autonomously use search engines to collect information through reinforcement learning. By incorporating this autonomous knowledge retrieval into the reasoning process, these systems are able to act with a more refined understanding and adapt their responses to changing tasks. Gemini 2.0 has the ability to process and generate multi-modal content. By deeply integrating with Google’s various tools and combining its advanced reasoning capabilities, it can effectively decompose tasks and gradually obtain the required information when dealing with multi-step problems. While current models have demonstrated initial versions of this functionality, they fall short in their ability to engage in sustained, interactive reasoning across diverse modalities.
@@
The behaviors of multimodal agents are closely linked to the deep reasoning capabilities of the underlying large multimodal models, particularly in terms of perception range, understanding accuracy, and reasoning depth. Thus, developing a comprehensive omni-modal model for real-world applications and enhancing its deep reasoning ability is foundational.
Early work, AnyGPT (Zhan et al., 2024), utilizes discrete representations for the unified processing of various modalities, achieving unified understanding and generation across modalities. Recently, Baichuan-Omni-1.5 (Li et al., 2025h) showcases impressive capabilities in collaborative real-time understanding across various modalities. Qwen2.5-Omni (Xu et al., 2025b) uses a new position embedding, named Time-aligned Multimodal RoPE, to synchronize the timestamps of video inputs with audio. More latest open source work, like M2-omni (Guo et al., 2025) and MiniCPM-o (Yu et al., 2024b), is narrowing the performance gap with closed-source models like GPT-4o.
Driven by real-world specific needs, omni-modal models with smaller size are gaining more and more attention. Megrez-3B-Omni (Li et al., 2025a) is an on-device multimodal understanding LLM model that demonstrates excellent performance in tasks such as scene understanding and OCR. Mini-Omni2 (Xie & Wu, 2024), a visual-audio assistant capable of providing real-time, end-to-end voice responses to visoin and audio queries. R1-Omni (Zhao et al., 2025c) focuses on emotion recognition from visual and auditory information.
@@
Figure 9: Overview of next-generation native large multimodal reasoning model. The envisioned system aims to achieve comprehensive perception across diverse real-world data modalities, enabling precise omnimodal understanding and in-depth generative reasoning. This foundational model will lead to more advanced forms of intelligent behavior, learning from world experience and realizing lifelong learning and self-improvement.
@@ 4.3 Technical Prospects
The technical prospect of Native Large Multimodal Reasoning Models (N-LMRMs) aims to natively unify understanding, generation, and reasoning across diverse data types, from language and vision to audio, tactile, sensor readings, temporal sequences, and structured data, bringing us closer to systems that can see, hear, talk, and act in a unified and cohesive manner. However, building such N-LMRMs poses significant challenges. These models must be architecturally designed to handle heterogeneous modalities within a single system, genetically use and combine diverse tools through long multimodal reasoning chains, and support continuous learning from real-world interactions. This section outlines key challenges in building N-LMRMs and proposes several potential pathways to address them.
Unified Representations and Cross-Modal Fusion.
@@
A fundamental challenge is creating a single model architecture that can process and generate different modalities in a coherent way. Traditional approaches often use separate encoders for each modality (Lyu et al., 2023; Li et al., 2024l). In contrast, native omni-modal models seek a more unified design that allows for seamless interaction between modalities. One possible solution is to homogenize all inputs and outputs into a common format and process any modality uniformly. This approach requires careful design to prevent negative interference, where one modality may dominate or impair the representation of others (Leng et al., 2024; Chen et al., 2024g). Thus an emerging solution is Mixture-of-Experts (MoE) architectures, with experts specialized for certain modalities are only activated for relevant inputs, while a core language model serves as the backbone for language intelligence (Chen et al., 2024i; Li et al., 2025j; Team, 2025a; Shukor et al., 2025).
Interleaved Multimodal Long Chain-of-Thought.
@@
Building on unified representations, N-LMRMs can extend traditional long internal chains of thought into interleaved reasoning processes across multiple modalities. This enables a new axis for test-time compute scaling that seamlessly blends different modalities (Wang et al., 2025a). OpenAI’s recently released o3 and o4-mini represent pioneering steps in this direction, i.e. reasoning with images in their chain of thought (OpenAI, 2025b), by automatically processing with tools that can zoom, crop, flip, or enhance images. Importantly, these capabilities come natively, without relying on separate specialized models (Wu & Xie, 2023; Hu et al., 2024b; Feng et al., 2025a; Qian et al., 2025; Wang et al., 2025d). Driven by the promising generalization capabilities of reinforcement learning across domains such as software engineering (OpenAI, 2025), IMO-level math (DeepSeek-AI et al., 2025), creative writing (Zhao et al., 2024c), and GUI manipulation (Qin et al., 2025), scaling reinforcement learning to more modalities, longer tool-augmented reasoning chains, and a broader set of reasoning tasks could be the recipe for the next generation of N-LMRMs, capable of simulating cross-modal reasoning and elevating machine intelligence.
Learning and Evolving from World Experiences.
@@
In dynamically evolving intelligent systems, the core value of LMRMs-based “World Model [^1] https://sites.google.com/view/worldmodel-iclr2025/” lies not only in its real-time modelling and reasoning capabilities in complex environments, like autonomous driving (Wang et al., 2024m) but also in its evolutionary mechanism for life-long learning (Thrun & Mitchell, 1995) through continuous interaction with the environment. When the MCP and A2A create a high-density network of tools and agent clusters, the system can transform each interaction into structured experiences through multidimensional engagement with the environment, tools, and other agents. This includes everything from pattern recognition in real-time data streams to causal reasoning across tool operation chains, from collaborative feedback in communication networks to autonomous adaptation in abnormal scenarios.
@@
This continuous learning paradigm enables LMRMs to overcome the limitations of static knowledge bases. By iteratively accumulating world experiences, it dynamically updates its cognitive architecture and decision-making strategies. Particularly in open environments, the autonomous learning mechanism drives the model to actively explore the potential of tool combinations. In the process of solving new problems, it simultaneously stores transferable knowledge, ultimately forming an intelligent system that possesses specialized reasoning capabilities while maintaining cross-scenario generalization resilience. We think the interactive learning method of online reinforcement learning and offline verification methods may iteratively and continuously stimulate the capabilities of LMRMs, which have been utilized in the GUI agentic model (Qin et al., 2025; Zheng et al., 2025a; Wang et al., 2024n) to continually improve the performance.
Data Synthesis.
@@
The current capabilities of LMRMs are largely data-driven. To enhance these models during the pre-training stage, it is crucial to develop a high-quality data synthesis pipeline that tailors their functionalities. Most existing efforts (Chang et al., 2024; Huang et al., 2025c; Xu et al., 2024c) in data synthesis focus on improving single-modal or cross-modal understanding and reasoning, particularly in domains like vision, language, and speech. However, there has been limited exploration of more complex aspects, such as aligning three or more modalities, creating multimodal interactive chains of thought and visual generation, implementing multi-step planning in dynamic environments, and coordinating multi-tool calls and parallel tool usage. These areas present significant opportunities for advancing multimodal reasoning models.
@@
In conclusion, we introduce the concept of N-LMRM as an initial step towards transitioning from capable reasoners to autonomous agents. Additionally, in alignment with OpenAI’s five-stage pathway to AGI (OpenAI, 2023), we are laying the groundwork for subsequent stages, including self-evolving innovators (Yamada et al., 2025) and multi-agent organizations (Zhang et al., 2025d). Building on our research proposal, future work can explore more agentic and omni-modal capabilities, advancing the development of increasingly autonomous machine intelligence.

### v2

@@ 3.3.3 Multimodal-R1
Tables 9-11 present a comprehensive survey of approximately 100 recent research papers that utilise reinforcement learning (RL) algorithms to enhance the long-chain reasoning capabilities of large multimodal models. Specifically, DPO in reinforcement learning has been widely used to enhance the reasoning capabilities of large multimodal models in recent years. RLHF-V (Yu et al., 2024a), LLaVA-Reasoner (Zhang et al., 2024e) and Insight-V (Dong et al., 2024b), by leveraging a large amount of self-constructed preference data and directly applying the DPO algorithm for training, have somewhat improved the reasoning ability of the models. MMPR (Wang et al., 2024g) made modifications to the DPO algorithm, adding quality loss obtained from a Binary Classifier and generation loss from traditional SFT on top of the DPO Preference loss, which effectively enhanced the model’s CoT capabilities.
With the success of Deepseek-R1, the GRPO algorithm began to be widely applied in multimodal large models. Works such as MM-EUREKA (Meng et al., 2025),Vt-R1 (Zhou et al., 2025b), LMM-R1 (Yingzhe et al., 2025), R1-V (Chen et al., 2025c) have adopted approaches similar to those in the text domain. They successfully applied the GRPO algorithm to mathematical geometry problems, demonstrating the phenomenon of reflection. Other works including VLM-R1 (Shen et al., 2025b), Visual-RFT (Liu et al., 2025l), ViCrit (Wang et al., 2025s) and Seg-Zero (Yuqi et al., 2025) utilize the GRPO algorithm to enhance the visual capabilities of multimodal large language models, such as grounding, detection, understanding and classification. Works such as SynthRL (Wu et al., 2025g) and MoDoMoDo (Liang et al., 2025c) improve GRPO performance through data augmentation, highlighting the critical role of high-quality reasoning data in RL. While most existing GRPO approaches focus on image-based tasks, some works have begun extending this algorithm into video and audio modalities. Video-R1 (Feng et al., 2025b) and VideoChat-R1 (Li et al., 2025j) introduce the GRPO algorithm into the video understanding, while R1-Omni (Zhao et al., 2025d) and AV-Reasoner (Lu et al., 2025b) has further extended it to the audio modality. Additionally, works such as R1-Reward (Zhang et al., 2025n), UnifiedReward-Think (Wang et al., 2025u) and Mixed-R1(Xu et al., 2025d) emphasizes the crucial role of rule-based reward design and reward model training, demonstrating significant improvements on training stability and performance through high-quality reward signals derived from carefully designed or powerful reward models. Despite these successes, existing work is often limited to specific tasks, and current multimodal large models have not yet been able to generalise the long-chain-of-thought abilities learned from tasks such as mathematics to the model’s general capabilities, as seen with Deepseek-R1.
@@ 4.1 Experimental Findings
EmbodiedBench (Yang et al., 2025c) Multimodal Understanding, Spatial Reasoning T, I High and Low Action Levels
EmbodiedEval (Cheng et al., 2025a) Attribute QA, Spatial Reasoning T, I Broad Abilities Assessment
@@
BrowseComp (Wei et al., 2025b) Web Browsing T Easy to Verify but Hard to Solve
SWE-Bench Multimodal (Yang et al., 2024c) Code T, I Image Included in Problem Statement
@@
OmniMMI (Wang et al., 2025v) VQA, Proactive Reasoning T, V, A In Streaming Video Context
@@
Recent studies have introduced a series of omni-modal benchmarks designed to evaluate the ability of LMRMs to perform unified understanding and reasoning across various data types (e.g. images, audio, text, and video). For example, OmniMMI (Wang et al., 2025v) aims to comprehensively assess the interactive capabilities of streaming video contexts in open-world environments. Experimental results reveal that even commercial models, such as Gemini-1.5-Pro and GPT-4o, achieve an average accuracy of less than 20%. When tasks require unified modality understanding (OmniBench (Li et al., 2024j), TaskAnything and JudgeAnything (Pu et al., 2025), MixEvalL-X (Ni et al., 2024b)), the performance of both open-source and closed-source models is significantly lower than under single-modal conditions. Specifically, in the Audio-Video Question Answering (AVQA) task, such as WorldSense (Hong et al., 2025), Claude 3.5 Sonnet only achieves an average accuracy of 35%, while the best-performing open-source model only achieves an accuracy of 25%. In the case of more challenging multimodal reasoning tasks, such as BabelBench (Wang et al., 2024i) and OmnixR (Chen et al., 2024e), the performance of all models declines sharply as the number of modalities increases. This suggests that models struggle to generate reasoning paths for image, video, and audio inputs compared to text inputs. These findings collectively highlight that current LMRMs are not yet capable of effectively processing omni-modal inputs.
@@
A diverse range of tasks highlights the complexity and breadth of multimodal agent evaluation settings. These include AgentBench’s multi-environment tasks (Liu et al., 2023b, 2024d), WorFBench’s intricate workflow planning scenarios (Qiao et al., 2024), OSWorld’s and AndroidWorld’s full operating system interactions (Xie et al., 2024a; Rawles et al., 2024), EmbodiedBench’s vision-based navigation and manipulation challenges (Yang et al., 2025c), VisualWebArena’s visually grounded web tasks (Koh et al., 2024), and GAIA’s open-ended, tool-augmented queries (Hu et al., 2023). Together, these benchmarks span a wide spectrum of task types and modalities (e.g., text and vision), encompassing both realistic and tool-augmented environments.
Regarding the performance of LMRMs on agent benchmarks, these models generally lead to current performance and have made notable progress (Team, 2024, 2025a; Yao et al., 2024b). However, even state-of-the-art models consistently fall short of human-level reliability and struggle with complex, open-ended tasks. Evaluations across benchmarks repeatedly expose common bottlenecks: models often fail at real-world grounding (Gou et al., 2025a; Zheng et al., 2024a), coherent long-horizon reasoning and planning (Qian et al., 2025), seamless integration with external tools (Wang et al., 2025j), and maintaining robustness across diverse modalities and domains (Chu et al., 2025). For example, in the BrowseComp benchmark (Wei et al., 2025b), GPT-4o achieves only 0.6% accuracy, rising to 1.9% with browsing tools, highlighting weak tool-interactive planning capability. OpenAI’s reasoning model o1 reaches 9.9% but still leaves significant room for improvement. Notably, OpenAI Deep Research, with targeted tuning for web search, completes 51.5% of tasks via autonomous iterative tool calling and reasoning. The experimental results highlight that current large reasoning models remain deficient in long-horizon reasoning and adaptive planning, which may require specific tuning and architectural enhancements to evolve into truly native agentic systems.
@@
Recently, OpenAI released o3 and o4-mini, providing full agentic access to ChatGPT tools and enabling models to "think with images" OpenAI (2025b). The integration of visual content enhances multimodal reasoning directly within the thought process. For example, in Figure 6, o3 demonstrates a clear task decomposition during an 8-minute and 13-second thought process. It effectively determines the best way to crop each sub-figure through trial and error, ultimately arriving at the correct solution.
Beyond visual reasoning, we evaluated o3’s capabilities in file processing, puzzle solving, location identification, and multimedia content creation. As illustrated in Figure 7 and 8, o3 exhibits strong performance in complex multimodal problem-solving by capturing and leveraging subtle clues in images. However, several challenges are identified: 1) Language knowledge can interfere with visual input. As the finger counting case shown in Figure 8, o3 mistakenly identifies the image as the standard raised hand emoji showing four fingers plus a thumb, despite the image clearly displaying six fingers. 2) OpenAI o3 struggles with input file handling and multimedia content generation. Due to tool constraints and the lack of Internet access in coding environments, file processing and multimedia creation often result in inaccuracies. In the resume information collection case in Figure 8, phone numbers parsed from resume PDFs can be incorrect, and o3 hallucinates candidates’ project experiences by reusing similar content. Additionally, in multimedia creation cases in Figure 7, the generated frames fail to adhere to the "red panda" instructions, and o3 is unable to support interleaved text-image generation. 3) OpenAI o3 may fabricate reasoning in its thought process. It occasionally "lies" about its reasoning, constructing incorrect rationales for potentially correct answers (e.g., the puzzle-solving case in Figure 7). This problem needs to be solved urgently, as it could lead to the model attempting to deceive users during the post-training process. In fact, it highlights that the model has not yet mastered the relevant thinking logic to solve the problem.
@@ 4.2 Capability of N-LMRMs
Search-o1 (Li et al., 2025h) 32B T T Training-Free Multi-Hop QA, Math Agentic Search-Augmented Reasoning
@@
Embodied-Reasoner (Zhang et al., 2025i) 7B T, I T, A Pretrain+SFT GUI navigation Image-Text Interleaved Long-Horizon Embodied Reasoning
@@
Baichuan-Omni-1.5 (Li et al., 2025k) 7B T, I, A, V T, A Pretrain+SFT VQA, OCR, ASR, Math, GeneralQA Leading Medical Image Understanding
@@
R1-Omni (Zhao et al., 2025d) 0.5B T, A, V T RL Emotion Recognition RL with Verifiable Reward
Janus-Pro (Chen et al., 2025g) 1B, 7B T, I T, I Pretrain+SFT Multimodal Understanding, Text-to-Image Decoupling Visual Encoding For Understanding and Generation
AnyGPT (Zhan et al., 2024) 7B T, I, A T, I, A Pretrain+SFT Multimodal-to-Text and Text-to-Multimodal Discrete Representations For Unified Processing
Uni-MoE (Li et al., 2025n) 13B, 20B, 22B, 37B T, I, A, V T Pretrain+SFT VQA, AQA Modality-Specific Encoders with Connectors for Unified Representation
Ovis-U1 (Wang et al., 2025f) 3B T, I T, I Pretrain+SFT Multimodal understanding, T2I, Image Editing Unified training from LLM and diffusion decoder with token refiner
ShapeLLM-Omni (Ye et al., 2025a) 7B 3D, I, T 3D, T Pretrain+SFT Text-to-3D, Image-to-3D, 3D understanding, 3D editing Uses 3D VQVAE to tokenize meshes for a unified autoregressive framework
Ming-Lite-Omni (AI et al., 2025) 2.8B T, I, A, V I, T, A Pretrain+SFT Multimodal understanding & generation MoE LLM with modality-specific routers
BAGEL (Deng et al., 2025a) 14B T, I, V I, T, V Pretrain+SFT+RL Multimodal understanding & generation Unified decoder-only MoT architecture
@@
Moreover, Search-o1 (Li et al., 2025h) utilizes external knowledge retrieval during the reasoning process to fill gaps in their understanding. R1-Searcher (Song et al., 2025b) and DeepResearcher (Zheng et al., 2025e) enhance their ability to autonomously use search engines to collect information through reinforcement learning. By incorporating this autonomous knowledge retrieval into the reasoning process, these systems are able to act with a more refined understanding and adapt their responses to changing tasks. Gemini 2.0 has the ability to process and generate multi-modal content. By deeply integrating with Google’s various tools and combining its advanced reasoning capabilities, it can effectively decompose tasks and gradually obtain the required information when dealing with multi-step problems. While current models have demonstrated initial versions of this functionality, they fall short in their ability to engage in sustained, interactive reasoning across diverse modalities.
@@
The behaviours of multimodal agents are closely linked to the deep reasoning capabilities of the underlying large multimodal models, particularly in terms of perception range, understanding accuracy, and reasoning depth. Thus, developing a comprehensive omni-modal model for real-world applications and enhancing its deep reasoning ability is foundational.
Early work, AnyGPT (Zhan et al., 2024) utilizes discrete representations for the unified processing of various modalities, achieving unified understanding and generation across modalities. Recently, Baichuan-Omni-1.5 (Li et al., 2025k) showcases impressive capabilities in collaborative real-time understanding across various modalities. Qwen2.5-Omni (Xu et al., 2025b) uses a new position embedding, named Time-aligned Multimodal RoPE, to synchronize the timestamps of video inputs with audio. More latest open-source work, like M2-omni (Guo et al., 2025) and MiniCPM-o (Yu et al., 2024b), is narrowing the performance gap with closed-source models like GPT-4o. Recent Ming-Omni (AI et al., 2025) achieves a new state-of-the-art performance on any-modality inputs.
Driven by real-world specific needs, omni-modal models with smaller sizes are gaining more and more attention. Megrez-3B-Omni (Li et al., 2025a) is an on-device multimodal understanding LLM model that demonstrates excellent performance in tasks such as scene understanding and OCR. Mini-Omni2 (Xie & Wu, 2024), a visual-audio assistant capable of providing real-time, end-to-end voice responses to vision and audio queries. R1-Omni (Zhao et al., 2025d) focuses on emotion recognition from visual and auditory information.
@@
Figure 9: Overview of next-generation native large multimodal reasoning model. The envisioned system aims to achieve comprehensive perception across diverse real-world data modalities, enabling precise omnimodal understanding and in-depth generative reasoning. This foundational model will lead to more advanced forms of intelligent behaviour, learning from world experience and realizing lifelong learning and self-improvement.
@@ 4.3 Technical Prospects
The technical prospect of Native Large Multimodal Reasoning Models (N-LMRMs) aims to natively unify understanding, generation, and reasoning across diverse data, from language and vision to audio, tactile, sensor readings, temporal sequences, and structured data, bringing us closer to systems that can see, hear, talk, and act in a unified and cohesive manner. However, building such N-LMRMs poses significant challenges. These models must be architecturally designed to handle heterogeneous modalities within a single system, genetically use and combine diverse tools through long multimodal reasoning chains, and support continuous learning from real-world interactions. This section outlines key challenges in building N-LMRMs and proposes several potential pathways to address them.
@@
Unified Representations and Cross-Modal Fusion. A fundamental challenge is creating a single model architecture that can process and generate different modalities in a coherent way. Traditional approaches often use separate encoders for each modality (Lyu et al., 2023; Li et al., 2024l). In contrast, native omni-modal models seek a more unified design that allows for seamless interaction between modalities. One possible solution is to homogenize all inputs and outputs into a common format and process any modality uniformly. This approach requires careful design to prevent negative interference, where one modality may dominate or impair the representation of others (Leng et al., 2024; Chen et al., 2024g). Thus, an emerging solution is Mixture-of-Experts (MoE) architectures, where experts specialized for certain modalities are only activated for relevant inputs, while a core language model serves as the backbone for language intelligence (Chen et al., 2024i; Li et al., 2025n; Team, 2025a; Shukor et al., 2025).
@@
Interleaved Multimodal Long Chain-of-Thought. Building on unified representations, N-LMRMs can extend traditional long internal chains of thought into interleaved reasoning processes across multiple modalities. This enables a new axis for test-time compute scaling that seamlessly blends different modalities (Wang et al., 2025b). OpenAI’s recently released o3 and o4-mini represent pioneering steps in this direction, i.e. reasoning with images in their chain of thought (OpenAI, 2025b), by automatically processing with tools that can zoom, crop, flip, or enhance images. Importantly, these capabilities come natively, without relying on separate specialized models (Wu & Xie, 2023; Hu et al., 2024b; Feng et al., 2025a; Qian et al., 2025; Wang et al., 2025j). Driven by the promising generalization capabilities of reinforcement learning across domains such as software engineering (OpenAI, 2025), IMO-level math (DeepSeek-AI et al., 2025), creative writing (Zhao et al., 2024c), and GUI manipulation (Qin et al., 2025), scaling reinforcement learning to more modalities, longer tool-augmented reasoning chains, and a broader set of reasoning tasks could be the recipe for the next generation of N-LMRMs, capable of simulating cross-modal reasoning and elevating machine intelligence.
@@
Learning and Evolving from World Experiences. In dynamically evolving intelligent systems, the core value of LMRMs-based “World Model [^1] https://sites.google.com/view/worldmodel-iclr2025/” lies not only in its real-time modelling and reasoning capabilities in complex environments, like autonomous driving (Wang et al., 2024m) but also in its evolutionary mechanism for life-long learning (Thrun & Mitchell, 1995) through continuous interaction with the environment. When the MCP and A2A create a high-density network of tools and agent clusters, the system can transform each interaction into structured experiences through multidimensional engagement with the environment, tools, and other agents. This includes everything from pattern recognition in real-time data streams to causal reasoning across tool operation chains, from collaborative feedback in communication networks to autonomous adaptation in abnormal scenarios.
@@
This continuous learning paradigm enables LMRMs to overcome the limitations of static knowledge bases. By iteratively accumulating world experiences, it dynamically updates its cognitive architecture and decision-making strategies. Particularly in open environments, the autonomous learning mechanism drives the model to actively explore the potential of tool combinations. In the process of solving new problems, it simultaneously stores transferable knowledge, ultimately forming an intelligent system that possesses specialized reasoning capabilities while maintaining cross-scenario generalization resilience. We think the interactive learning method of online reinforcement learning and offline verification methods may iteratively and continuously stimulate the capabilities of LMRMs, which have been utilized in the GUI agentic model (Qin et al., 2025; Zheng et al., 2025a; Wang et al., 2024o) to continually improve the performance.
@@
Data Synthesis. The current capabilities of LMRMs are largely data-driven. To enhance these models during the pre-training stage, it is crucial to develop a high-quality data synthesis pipeline that tailors their functionalities. Most existing efforts (Chang et al., 2024; Huang et al., 2025d; Xu et al., 2024c) in data synthesis focus on improving single-modal or cross-modal understanding and reasoning, particularly in domains like vision, language, and speech. However, there has been limited exploration of more complex aspects, such as aligning three or more modalities, creating multimodal interactive chains of thought and visual generation, implementing multi-step planning in dynamic environments, and coordinating multi-tool calls and parallel tool usage. These areas present significant opportunities for advancing multimodal reasoning models.
@@
In conclusion, we introduce the concept of N-LMRM as an initial step towards transitioning from capable reasoners to autonomous agents. Additionally, in alignment with OpenAI’s five-stage pathway to AGI (OpenAI, 2023), we are laying the groundwork for subsequent stages, including self-evolving innovators (Yamada et al., 2025) and multi-agent organizations (Zhang et al., 2025l). Building on our research proposal, future work can explore more agentic and omni-modal capabilities, advancing the development of increasingly autonomous machine intelligence.